# ResearchArena: Benchmarking Large Language Models’ Ability to Collect and Organize Information as Research Agents

### v2

@@ Abstract
Large language models (LLMs) excel across many natural language processing tasks but face challenges in domain-specific, analytical tasks such as conducting research surveys. This study introduces ResearchArena, a benchmark designed to evaluate LLMs’ capabilities in conducting academic surveys—a foundational step in academic research. ResearchArena models the process in three stages: (1) information discovery, identifying relevant literature; (2) information selection, evaluating papers’ relevance and impact; and (3) information organization, structuring knowledge into hierarchical frameworks such as mind-maps. Notably, mind-map construction is treated as a bonus task, reflecting its supplementary role in survey-writing. To support these evaluations, we construct an offline environment of 12M full-text academic papers and 7.9K survey papers. To ensure ethical compliance, we do not redistribute copyrighted materials; instead, we provide code to construct the environment from the Semantic Scholar Open Research Corpus (S2ORC). Preliminary evaluations reveal that LLM-based approaches underperform compared to simpler keyword-based retrieval methods, underscoring significant opportunities for advancing LLMs in autonomous research.
@@ 1 Introduction
Preliminary evaluations reveal that LLMs underperform compared to simpler keyword-based search methods in tasks requiring analytical depth. For example, using survey titles as retrieval queries consistently yields superior recall and precision compared to LLM-driven information discovery and selection tasks. Additionally, under the task of information organization, LLMs face challenges in constructing coherent structures without the oracle guidance, underscoring the need for improvements in organizational and analytical capabilities.
The dataset supporting ResearchArena comprises 12M full-text academic papers and 7.9K survey papers, curated from the Semantic Scholar Open Research Corpus (S2ORC) Lo et al. (2019). This corpus ensures scholarly rigor and relevance, offering a robust foundation for benchmarking LLM performance across diverse domains. Furthermore, our released pipeline supports weekly updates from Semantic Scholar, which enables evaluations to incorporate recent advancements, ensuring ongoing relevance.
@@ 4 Analysis
Disciplinary Distribution. We classified each of the 12.0M papers in our public corpus and 7.9K survey papers by the top-5 most popular academic disciplines. This classification was based on the indexing information provided by S2ORC. Frequencies of papers per discipline were then aggregated and visualized to identify trends and imbalances. Figure 2(a) and 2(b) revealed significant disparities in the frequency of disciplines between the public corpus and the survey subset. Notably, Computer Science is the most prevalent discipline within surveys but less common in the broader corpus. This could reflect the dynamic nature of the CS field, which often necessitates comprehensive reviews to synthesize rapid advancements and emerging trends.
@@ 5 Benchmark Tasks
For evaluation, two primary metrics are employed: Heading Soft Recall Fränti and Mariescu-Istodor (2023) and Heading Entity Recall Shao et al. (2024). These metrics compare the set of node labels from the original and the constructed knowledge mind-maps, referred to as $A$ and $B$, respectively. To measure similarity of these labels, Heading Soft Recall (HSR) leverages Sentence-Bert Reimers and Gurevych (2019) embedding, while Heading Entity Recall (HER) employs Named Entity Recognition from FLAIR Akbik et al. (2019) for extraction. The formal definitions for each metric are presented in Equation 1, where $S$ is the set of labels extracted from the mind-maps.
@@
$\text{C}(S)$ $=\sum_{i=1}^{|S|}\frac{1}{\sum_{j=1}^{|S|}\text{Sim}(S_{i},S_{j})}$ (1)
$\text{HSR}(A,B)$ $=\frac{\text{C}(A)+\text{C}(B)-\text{C}(A\cup B)}{\text{C}(B)}$
$\text{HER}(A,B)$ $=\frac{|\text{Ent}(A)\cap\text{Ent}(B)|}{|\text{Ent}(A)|}$
@@ 6 Benchmarking
Zero-Shot 0.0382 0.0832 0.1253 0.2287 0.0602 0.1232 0.0256 0.0464
Decomposer 0.0434 0.0879 0.1431 0.2554 0.0717 0.1304 0.0312 0.0536
@@
Zero-Shot 0.0634 0.1346 0.0692 0.1417 0.0856 0.1657 0.1743 0.3246
Decomposer 0.0735 0.1445 0.0803 0.1554 0.0986 0.1838 0.1959 0.3510
@@
In this section, we present preliminary evaluations of existing techniques, describing their configurations and performance metrics. These techniques encompass both naive keyword-based methods, such as Title, and advanced LLM-based methods, including STORM. The exact wording of the prompts used in each baseline can be found in Appendix B.
6.1 Baselines
  • Title: Assuming that research topics are encapsulated within survey titles, this method directly employs the title from each survey paper as a query to retrieve relevant materials that support research on the topic. It is important to note that title extraction using S2ORC exhibits variable capitalization across different documents. As a result, we normalize by converting titles to lowercase.
  • Zero-Shot: Assuming that existing LLMs possess prior knowledge relevant to a survey topic, this method extends the Title method by instructing GPT-4 to derive a query from the survey title. This approach leverages the inherent capabilities of LLMs to generate more sophisticated and contextually appropriate queries.
  • Decomposer: As discovered by Tushar et al. Khot et al. (2022), decomposed prompting is more effective when individual reasoning steps of a task are difficult to learn. This principle is applicable to our case, as a survey topic may consist of multiple sub-topics, making it challenging to directly generate a single query that retrieves all relevant papers. Consequently, we instruct GPT-4 to first deconstruct the research topic into several sub-questions. Each sub-question then generates a corresponding sub-query. These sub-queries are retrieved in batches, and the results are amalgamated using reciprocal rank fusion Cormack et al. (2009).
  • Self-RAG: As proposed by Asai et al. Asai et al. (2023), Self-RAG adaptively retrieves passages on demand and utilizes reflection tokens to determine which retrieved documents are relevant to the instruction, thus continuing the generation based on the pertinent information. It serves as an enhanced version of Zero-Shot, where the model is instructed to generate a query from the topic. Because the model refines its final query generation based on the discovered information from intermediate retrievals, it operates as a research agent.
  • STORM: As presented in Section 2, STORM conducts research through multi-perspective conversations to compose Wikipedia articles on particular topics from scratch. It closely resembles our scenario, except that the environment involves more rigorous academic papers. We record the retrieval history as STORM continues to probe for additional papers. Upon concluding the final round of conversations, every article within the retrieval history is considered part of the discovered information.
  • Deep Research: While Deep Research demonstrates advanced multi-step web-based browsing and iterative information synthesis, it does not currently support offline corpora and relies on general web pages, prompting us to include a brief case study of its capabilities.
6.2 Evaluation Results
@@
Deep Research. We conducted a brief investigation on two distinct topics, transfer learning and LiDAR scanning mechanisms, using Gemini Deep Research, as shown in Appendix D. Due to its sole reliance with online resources, we don’t have a direct quantitative comparison with other baselines. Nevertheless, the generated summaries illustrate Gemini Deep Research’s capacity to synthesize diverse online sources into coherent findings, highlighting its potential to support high-level exploration when specialized academic databases are not immediately required.
@@ 8 Limitations
Despite the robust framework and extensive dataset provided by ResearchArena, this study has several limitations. Firstly, the offline environment, though comprehensive, may not accurately represent the dynamic and interconnected nature of live databases and the internet. This discrepancy could potentially limit the applicability of the findings in real-world research settings. Additionally, due to copyright constraints, not every full-text reference of the survey papers could be included. This omission could affect the comprehensive understanding of the survey topics under investigation. Finally, there is no evaluation on text generation but mostly the surveying process. However, even if this is just the first step of conducting research, LLM agents have already shown deficiencies. Future iterations of ResearchArena should address this issue, particularly as these agents improve.
@@ Appendix C Inspection over Survey Passages
A manual inspection of survey passages is conducted to assess the effectiveness of GPT-4 in identifying survey papers. Following this, the extracted mind-map is analyzed and compared against the original paper to determine its accuracy and relevance in capturing key aspects of the topic under discussion.
CorpusID IsSurvey AccurateMap RelevantMap
3524264 True True True
3644401 True True True
4503761 True True False
5058972 True True False
8034133 True True True
8922493 True True True
9935621 True False True
10934716 True True True
17513321 True True True
18750590 True True True
20774863 False False False
22727391 True True True
23384543 True True True
33413610 True True False
52841074 True True True
55254540 True True True
56895323 True True True
57721147 True False True
64256250 True False True
115523285 True True False
119297355 True False True
198933686 True True True
210865204 True True False
211010433 True True True
212665971 True True True
214743520 True True True
216056393 True True True
219316962 True True False
220302470 True False True
221446014 True True True
222095837 True True False
225029039 False False False
226227376 True True False
226300094 True True True
227228021 True True False
229363354 True False True
231149960 False False False
231698518 True True True
233722066 False False False
234213205 True True False
235352671 True False False
235458292 True True True
235485414 True True False
235490196 True True True
235669589 True True True
235766219 True True False
236090307 True True True
236772630 True True True
236976256 True True True
236986986 True True False
CorpusID IsSurvey AccurateMap RelevantMap
237372527 True False True
237373628 True True True
237485263 True True True
238242214 True False True
238242941 True True True
238242941 True False False
238639787 True True False
244119139 True True True
244773222 True True True
245353469 True True True
245877584 False False False
247763152 True True True
248834382 True True False
249209981 True True False
249687282 True False False
250089226 True True True
250939903 True True True
251104722 True True True
251506514 True True True
251643467 True True False
252683270 True True True
252762319 True True False
253370610 True True True
253796900 True True True
254274880 True False True
254756520 True True True
255025269 True True True
256227178 True True True
256826729 True True True
257232619 True True False
257255597 True False True
257522478 True True False
258539255 True True True
258722762 True True True
259043594 True True True
259088696 True True True
259108865 True True True
259154569 False False False
259283502 True True True
259951356 True True True
260229544 True True True
260316174 True True True
260849783 True False True
261682162 True False True
263134374 True True False
263334211 True True True
263830273 True True True
263831409 True False True
263909496 True True True
264604532 True True True
Appendix D Case Study with Gemini Deep Research

### v3

@@ Abstract
Large language models (LLMs) excel across many natural language processing tasks but face challenges in domain-specific, analytical tasks such as conducting research surveys. This study introduces ResearchArena, a benchmark designed to evaluate LLMs’ capabilities in conducting academic surveys—a foundational step in academic research. ResearchArena models the process in three stages: (1) information discovery, identifying relevant literature; (2) information selection, evaluating papers’ relevance and impact; and (3) information organization, structuring knowledge into hierarchical frameworks such as mind-maps. Notably, mind-map construction is treated as a bonus task, reflecting its supplementary role in survey-writing. To support these evaluations, we construct an offline environment of 12M full-text academic papers and 7.9K survey papers. To ensure ethical compliance, we do not redistribute copyrighted materials; instead, we provide code to construct the environment from the Semantic Scholar Open Research Corpus (S2ORC). Preliminary evaluations reveal that LLM-based approaches underperform compared to simpler keyword-based retrieval methods, though recent reasoning models such as DeepSeek-R1 show slightly better zero-shot performance. These results underscore significant opportunities for advancing LLMs in autonomous research. We open-source the code to construct the ResearchArena benchmark at https://github.com/cxcscmu/ResearchArena.
@@ 1 Introduction
Preliminary evaluations reveal that LLMs underperform compared to simpler keyword-based search methods in tasks requiring analytical depth. For example, using survey titles as retrieval queries consistently yields superior recall and precision compared to LLM-driven information discovery and selection tasks. Additionally, under the task of information organization, LLMs face challenges in constructing coherent structures without the oracle guidance, underscoring the need for improvements in organizational and analytical capabilities. Nonetheless, we observe promising results from reasoning-oriented models such as DeepSeek-R1 (Guo et al., 2025), which show modest gains in zero-shot performance—suggesting that improved reasoning capabilities may help bridge some of these gaps.
The dataset supporting ResearchArena comprises 12M full-text academic papers and 7.9K survey papers, curated from the Semantic Scholar Open Research Corpus (S2ORC) Lo et al. (2019). This corpus ensures scholarly rigor and relevance, offering a robust foundation for benchmarking LLM performance across diverse domains.
@@ 3 Collection Methodology
Figure 1: Schematic overview of the construction pipeline for ResearchArena.
@@ 3.4 Dataset Access
Most importantly, our released pipeline supports weekly updates from Semantic Scholar, enabling evaluations to continuously incorporate recent advancements and maintain ongoing relevance.
@@ 4 Analysis
Disciplinary Distribution. We classified each of the 12.0M papers in our public corpus and 7.9K survey papers by the top-5 most popular academic disciplines. This classification was based on the indexing information provided by S2ORC. Frequencies of papers per discipline were then aggregated and visualized to identify trends and imbalances. Figure 3(a) and 3(b) revealed significant disparities in the frequency of disciplines between the public corpus and the survey subset. Notably, Computer Science is the most prevalent discipline within surveys but less common in the broader corpus. This could reflect the dynamic nature of the CS field, which often necessitates comprehensive reviews to synthesize rapid advancements and emerging trends. On the other hand, the dominance of medical papers in the public corpus stems from the composition of the original S2ORC dataset, which was itself heavily sourced from biomedical publications.
@@ 5 Benchmark Tasks
For evaluation, two primary metrics are employed: Heading Soft Recall Fränti and Mariescu-Istodor (2023) and Heading Entity Recall Shao et al. (2024). These metrics compare the set of node labels from the original and the constructed knowledge mind-maps, referred to as $A$ and $B$, respectively. To measure similarity of these labels, Heading Soft Recall (HSR) leverages Sentence-Bert Reimers and Gurevych (2019) embedding, while Heading Entity Recall (HER) employs Named Entity Recognition from FLAIR Akbik et al. (2019) for extraction.
@@ 6 Benchmarking
Zero-Shot (GPT-4) 0.0382 0.0832 0.1253 0.2287 0.0602 0.1232 0.0256 0.0464
Decomposer (GPT-4) 0.0434 0.0879 0.1431 0.2554 0.0717 0.1304 0.0312 0.0536
Zero-Shot (Claude 3.5 Sonnet) 0.0336 0.0777 0.1173 0.2169 0.0537 0.1137 0.0235 0.0428
Decomposer (Claude 3.5 Sonnet) 0.0435 0.0876 0.1496 0.2547 0.0751 0.1290 0.0331 0.0541
Zero-Shot (DeepSeek-R1) 0.0446 0.0865 0.1491 0.2459 0.0730 0.1304 0.0316 0.0511
Decomposer (DeekSeek-R1) 0.0418 0.0858 0.1478 0.2514 0.0731 0.1281 0.0331 0.0541
@@
Zero-Shot (GPT-4) 0.0634 0.1346 0.0692 0.1417 0.0856 0.1657 0.1743 0.3246
Decomposer (GPT-4) 0.0735 0.1445 0.0803 0.1554 0.0986 0.1838 0.1959 0.3510
Zero-Shot (Claude 3.5 Sonnet) 0.0584 0.1258 0.0630 0.1326 0.0792 0.1562 0.1763 0.3097
Decomposer (Claude 3.5 Sonnet) 0.0760 0.1437 0.0827 0.1548 0.1025 0.1834 0.2015 0.3518
Zero-Shot (DeepSeek-R1) 0.0756 0.1401 0.0824 0.1488 0.1016 0.1763 0.2081 0.3336
Decomposer (DeekSeek-R1) 0.0740 0.1424 0.0811 0.1530 0.1009 0.1817 0.1977 0.3520
@@
In this section, we present preliminary evaluations of existing techniques, describing their configurations and performance metrics. These techniques encompass both naive keyword-based methods, such as Title, and advanced LLM-based methods, including STORM. The exact wording of the prompts used in each baseline can be found in Appendix B. We adopt the abstract as the indexing scheme, as it consistently yields the best performance. Results for alternative indexing schemes are reported in Appendix D.
6.1 Baselines
Title: This method assumes that survey titles encapsulate their respective research topics. It directly uses the title from each survey paper as a query to retrieve relevant materials. Since titles in the S2ORC corpus exhibit inconsistent capitalization, we normalize them by converting all titles to lowercase.
Zero-Shot: Building on the Title method, this approach leverages LLMs—such as GPT-4, Claude 3.5 Sonnet, or DeepSeek-R1—to generate refined queries based on the survey title. The assumption is that these models possess prior knowledge relevant to the topic and can produce more precise and context-aware queries.
Decomposer: Based on the findings of Tushar et al. Khot et al. (2022), decomposed prompting is beneficial when complex reasoning tasks are difficult to perform in a single step. Survey topics often consist of multiple sub-topics, making them well-suited for this approach. We instruct LLMs to break the topic into several sub-questions, each of which is used to generate a sub-query. These sub-queries are executed in batches, and the retrieved results are merged using reciprocal rank fusion Cormack et al. (2009).
Self-RAG: Proposed by Asai et al. Asai et al. (2023), Self-RAG retrieves passages on demand and uses reflection tokens to assess their relevance. It then incorporates relevant information into the ongoing generation. This approach improves upon Zero-Shot by allowing the model to iteratively refine its queries based on intermediate retrievals, effectively acting as a research agent.
STORM: As described in Section 2, STORM uses multi-perspective conversations to construct Wikipedia-style articles from scratch. It mirrors our use case, except that our context involves academic literature rather than encyclopedic content. We log all retrieved documents during the conversation process, and the final retrieval set is treated as the body of discovered evidence.
Deep Research: This system showcases advanced multi-step browsing and iterative information synthesis via web-based exploration. However, it currently does not support offline corpora and relies on open web resources. We include a brief case study to illustrate its potential and limitations.
6.2 Evaluation Results
@@
Deep Research. We conducted a brief investigation on two distinct topics, transfer learning and LiDAR scanning mechanisms, using Gemini Deep Research, as shown in Appendix E. Due to its sole reliance with online resources, we don’t have a direct quantitative comparison with other baselines. Nevertheless, the generated summaries illustrate Gemini Deep Research’s capacity to synthesize diverse online sources into coherent findings, highlighting its potential to support high-level exploration when specialized academic databases are not immediately required.
@@ 8 Limitations
Despite the robust framework and extensive dataset provided by ResearchArena, this study has several limitations. Firstly, due to copyright constraints, not every full-text reference of the survey papers could be included. This omission could affect the comprehensive understanding of the survey topics under investigation. In addition, there is no evaluation on text generation but mostly the surveying process. However, even if this is just the first step of conducting research, LLM agents have already shown deficiencies. Future iterations of ResearchArena should address this issue, particularly as these agents improve.
@@ Appendix C Inspection over Survey Passages
A manual inspection of survey passages is conducted to assess the effectiveness of GPT-4 in identifying survey papers. Following this, the extracted mind-map is analyzed and compared against the original paper to determine its accuracy and relevance in capturing key aspects of the topic under discussion. Details are presented in Table 6 and 7.
Appendix D Indexing Schemes
We present a comparative analysis of different document indexing schemes—title-only, abstract, and full-text—used during the information discovery stage. While the main experiments adopt abstract indexing for its balance between informativeness and tractability, we evaluated how alternative schemes affect retrieval performance across three baselines: Title, Zero-Shot, and Decomposer using GPT-4. Results are reported for both BM25 and BGE retrieval models in Table 5.
Table 5: Impact of indexing schemes on retrieval performance.
Recall@10 Recall@100 Precision@10 Precision@100
Method (Indexing) BM25 BGE BM25 BGE BM25 BGE BM25 BGE
Title (Title) 0.0150 0.0184 0.0347 0.0419 0.0297 0.0378 0.0082 0.0101
Title (Full-Text) 0.0038 0.0018 0.0072 0.0027 0.0091 0.0054 0.0020 0.0009
Zero-Shot (Title) 0.0122 0.0122 0.0292 0.0293 0.0245 0.0238 0.0069 0.0067
Zero-Shot (Full-Text) 0.0035 0.0016 0.0067 0.0024 0.0081 0.0046 0.0019 0.0008
Decomposer (Title) 0.0125 0.0130 0.0297 0.0305 0.0238 0.0250 0.0070 0.0072
Decomposer (Full-Text) 0.0037 0.0017 0.0076 0.0026 0.0088 0.0049 0.0022 0.0008
Appendix E Case Study with Gemini Deep Research
@@
CorpusID IsSurvey AccurateMap RelevantMap
3524264 True True True
3644401 True True True
4503761 True True False
5058972 True True False
8034133 True True True
8922493 True True True
9935621 True False True
10934716 True True True
17513321 True True True
18750590 True True True
20774863 False False False
22727391 True True True
23384543 True True True
33413610 True True False
52841074 True True True
55254540 True True True
56895323 True True True
57721147 True False True
64256250 True False True
115523285 True True False
119297355 True False True
198933686 True True True
210865204 True True False
211010433 True True True
212665971 True True True
214743520 True True True
216056393 True True True
219316962 True True False
220302470 True False True
221446014 True True True
222095837 True True False
225029039 False False False
226227376 True True False
226300094 True True True
227228021 True True False
229363354 True False True
231149960 False False False
231698518 True True True
233722066 False False False
234213205 True True False
235352671 True False False
235458292 True True True
235485414 True True False
235490196 True True True
235669589 True True True
235766219 True True False
236090307 True True True
236772630 True True True
236976256 True True True
236986986 True True False
CorpusID IsSurvey AccurateMap RelevantMap
237372527 True False True
237373628 True True True
237485263 True True True
238242214 True False True
238242941 True True True
238242941 True False False
238639787 True True False
244119139 True True True
244773222 True True True
245353469 True True True
245877584 False False False
247763152 True True True
248834382 True True False
249209981 True True False
249687282 True False False
250089226 True True True
250939903 True True True
251104722 True True True
251506514 True True True
251643467 True True False
252683270 True True True
252762319 True True False
253370610 True True True
253796900 True True True
254274880 True False True
254756520 True True True
255025269 True True True
256227178 True True True
256826729 True True True
257232619 True True False
257255597 True False True
257522478 True True False
258539255 True True True
258722762 True True True
259043594 True True True
259088696 True True True
259108865 True True True
259154569 False False False
259283502 True True True
259951356 True True True
260229544 True True True
260316174 True True True
260849783 True False True
261682162 True False True
263134374 True True False
263334211 True True True
263830273 True True True
263831409 True False True
263909496 True True True
264604532 True True True